# Reinforced Generation of Combinatorial Structures: Hardness of Approximation

### v5

@@ 1 Introduction
    Hardness of certifying upper bounds on the MAX-CUT and MAX-IND-SET of sparse random graphs. In Section 2, we improve upon the combinatorial structures presented in [36], achieving improved lower bounds in the $\{3,4\}$-regular cases. Complementing this result, we obtain new upper bounds which improve upon Hoffman’s classic spectral bounds [30, 25]. (These upper bounds are derived via analytical approaches [^1] These improvements are derived from a discussion with Sidhanth Mohanty, Northwestern University..) Together, these achieve an almost tight resolution to the problem. [^2] By this we mean: the lower bounds and upper bounds on efficient certification algorithms match up to the second decimal place..
@@
    NP-Hardness of approximation for metric TSP. In Section 4, we give a new gadget based reduction (from standard Håstad PCPs [26]) for the NP-hardness of approximating metric TSP. We improve the current SOTA of $117/116$ [11] to $111/110$. Our improvement advances the long line of work establishing hardness results for this problem [45, $1+\delta$] (without an explicit value for $\delta$), [16, 5381/53805381/5380][7, 3183/31823183/3182][44, 220/219220/219][37, 185/184185/184][33, 123/122123/122][11, $117/116$]. As an aside, our AI-based exploration is ongoing, with the potential for further improvements to this bound.
Notably, our results come from a single (AI-derived) technique, AlphaEvolve [42, 48], to find and verify finite combinatorial structures that improve on prior constructions. At a high level AlphaEvolve uses an LLM (Large-Language-Model) to iteratively evolve code snippets that generate combinatorial structures (we will sometimes refer to these as constructions) that are of high quality by some criterion. While the structures we generate using AlphaEvolve are finite constructions, our main results (Theorems 2.13.1, and 4.1) embed the universal quantifier $\forall n$, with $n$ being the size of the problem instance, thanks to appropriate “lifting” [36, 50, 11] arguments. At a minimum our work suggests that we would do well to routinely run gadget proofs through an “AI optimization” phase. We present our results in increasing order of human involvement (required to make the problem amenable to AlphaEvolve).
@@
For the hardness of certifying upper bounds for sparse random graphs, we use AlphaEvolve to search for $d=\{3,4\}$-regular Ramanujan graphs, along with a witness for a large cut/independent set. Evaluating whether a graph is Ramanujan, and whether or not the cut/independent set is a valid witness, are efficient. As a result, for these problems we could find and verify graphs with as many as $163$ vertices.
@@
We emphasize that our theorems do not depend on the correctness of these improved verifiers; even though our AlphaEvolve computations internally use them, the gadgets found by AlphaEvolve were verified by a brute-force algorithm that explicitly checks all $\Omega(k^{m})$ constraints.
2 Average Case Hardness for Random Graphs
A central problem in average case complexity is that of certification, where our goal is to certify a property of inputs from some distribution $\mu=\mu_{n}$ over a sample space $\Omega=\Omega_{n}$. In particular, let $P:\Omega\to\{0,1\}$ be a property such that $\mathop{\mathbf{Pr}}_{x\sim\mu}[P(x)=1]=1-o_{n}(1)$. We are given as input a sample $x\sim\mu$, and the goal is to efficiently certify that $x$ satisfies $P$ with high probability.
@@ 2 Average Case Hardness for Random Graphs
Here, $\mathcal{G}(n,d)$ is the uniform distribution over random $d$-regular graphs for some fixed constant $d\in\mathbb{N}$.
We focus specifically on the setting of refutation, where $P$ is a co-NP property. This setting has seen considerable activity, leading to innovation in various algorithmic techniques such as spectral and sum-of-squares algorithms, along with the development of a variety of methods to establish hardness (for instance, [8, 49, 6, 36, 35]).
In particular, we will consider the problems of certifying properties of sparse random graphs sampled from $\mathcal{G}(n,d)$. The main question we consider is the following: what is the strongest upper bound one can efficiently certify on the MAX-CUT (or MAX-IND-SET) of a random $d$-regular graph?
@@
There is evidence that this is a nontrivial problem exhibiting a certification gap, in the sense that $\textup{MC}(G)$ concentrates around a constant value strictly smaller than $\sigma^{\textup{MC}}_{d}$ for $G\sim\mathcal{G}(n,d)$ [6, 36]. So in some sense, these are the simplest refutation-style properties that exhibit interesting complexity-theoretic phenomena. In this work, we focus on accurately characterizing the value of $\sigma^{\textup{MC}}_{d}$ and $\sigma^{\textup{IS}}_{d}$. We remark that the asymptotics of these quantities as $d\to\infty$ are already understood (thanks to [6, 36]). In service of a precise understanding of this problem in all regimes, we focus on the least understood case where $d$ is small; specifically we study $d\in\{3,4\}$.
Towards this goal, [36] introduced a conjecture [^3] We refer the reader to [36] for supporting evidence and intuition. that relates $\sigma^{\textup{MC}}_{d}$ and $\sigma^{\textup{IS}}_{d}$ to a fundamental question in spectral graph theory, specifically about Ramanujan graphs, which are (in a spectral sense) the most “random” looking regular graphs constructed via a deterministic process.
@@
In other words, the existence of Ramanujan graphs with notably large MAX-CUT and MAX-IND-SET values implies computational hardness of our certification problem. In particular, [36] exhibited finite $d$-regular Ramanujan graphs for $d\in\{3,4\}$, yielding concrete lower bounds on $\sigma_{d}^{\textup{MC}}$ and $\sigma_{d}^{\textup{IS}}$ that are strong enough to prove a certification gap. On the other hand, there was a large gap between the best known upper bound [30, 25] and their lower bounds (see Table 1 for a comparison). Our results make progress towards a more precise understanding of this question. Firstly, we use AlphaEvolve to construct better $d$-regular Ramanujan graphs, exhibiting stronger lower bounds.
@@ 3 Gadget based NP-Hardness for Approximating MAX-k-CUT
Let $\Psi$ be a finite alphabet, and let $\Gamma_{1},\ldots,\Gamma_{r}$ be a collection of predicates over $\Psi$, that is, $\Gamma_{i}:\Psi^{\ell_{i}}\to\{0,1\}$ for some $\ell_{i}\in\mathbb{N}$. We define $\textup{CSP}(\Gamma_{1},\ldots,\Gamma_{r})$ to be the collection of instances $I:\Psi^{n}\to\mathbb{R}_{\geq 0}$ of the form $I(\mathbf{x})=\sum_{j\in[m]}\zeta_{j}(\mathbf{x})$ where for each $j$, $\zeta_{j}$ is a clause, i.e., it takes the form $\zeta_{j}(\mathbf{x})=\Gamma_{i}(x_{a_{1}},x_{a_{2}},\ldots,x_{a_{\ell_{i}}})$ for some $i\in[r]$ and distinct $a_{1},\ldots,a_{\ell_{i}}\in[n]$.
In the Max-CSP setting, the input is a description of an instance $I$ of a particular CSP, and the goal is to compute the maximum number of constraints that can be satisfied. In other words, we are trying to approximately compute the function $f(I)=\max_{\mathbf{x}\in\Psi^{n}}I(\mathbf{x})$. It will be useful to cast the approximation problem as a decision problem, as we do below by introducing the notion of $(c,s)$-approximation.
@@
Given a constraint satisfaction problem $\textup{CSP}(\Gamma_{1},\ldots,\Gamma_{r})$ and constants $0\leq s\leq c\leq 1$, the $(c,s)$ approximation problem for $\textup{CSP}(\Gamma_{1},\ldots,\Gamma_{r})$ is the problem of, given an $n$-variable instance $I=\sum_{j\in[m]}\zeta_{j}$, deciding whether (1) $\max_{\mathbf{x}}I(\mathbf{x})\geq c\cdot m$ or (2) $\max_{\mathbf{x}}I(\mathbf{x})\leq s\cdot m$. Similarly for $0<\alpha\leq 1$, the $\alpha$-approximation problem is the problem of finding some $\mathbf{x}^{\prime}$ such that $I(\mathbf{x}^{\prime})\geq\alpha\cdot\max_{x}I(\mathbf{x})$.
@@
We have a strong quantitative understanding for a number of CSPs [26, 12] under NP-hardness. However, there are some quite simple CSPs for which very little is known quantitatively. For example, consider $\textup{MAX-}{k}\textup{-CUT}$, the problem of partitioning the vertices of a graph into $k$ sets in a way that maximizes the number of edges crossing the partition. This problem can be written as a CSP over alphabet $\Psi=\mathbb{Z}_{k}$ as $\textup{MAX-}{k}\textup{-CUT}:=\textup{CSP}(P^{\neq})$ where $P^{\neq}$ is the “inequality” predicate $P^{\neq}(x,y)=\mathbf{1}\{x\neq y\}$. For small values of $k$, there is a large gap between the best known algorithmic results (concretely, $0.836$ and $0.857$-approximations for $k=3$ and $k=4$ respectively [23, 14]) and the best known NP-hardness ($16/17+\varepsilon\approx 0.941$ and $85/86+\varepsilon\approx 0.9883$ for $k=3$ and $k=4$ respectively [3]) [^4] Even under the Unique Games Conjecture, the optimal approximability curve is not known for $k=4$, and for $k=3$ is only known in the low completeness regime [27] (which implies a weak inapproximability of $0.989$)..
@@
We use AlphaEvolve to find such gadget reductions from the well-understood CSP $3\textup{LIN}({k})$ to $\textup{MAX-}{k}\textup{-CUT}$ for $k\in\{3,4\}$. Here $3\textup{LIN}({k})$ is defined as $\textup{CSP}(P_{0}^{=},P_{1}^{=},\ldots,P_{k-1}^{=})$ for the $3$-ary “linear equation” predicates $P_{i}^{=}(x,y,z)=\mathbf{1}\{x+y+z\equiv i\pmod{k}\}$.
@@
For any $\varepsilon>0$, there is a gadget reduction from $(1-O(\varepsilon),1/3+O(\varepsilon))$-approximating $3\textup{LIN}({3})$ to $(0.9193-\varepsilon,0.887+\varepsilon)$-approximating $\textup{MAX-}{3}\textup{-CUT}$. As a consequence, it is NP-hard to $0.9649+\varepsilon$-approximate $\textup{MAX-}{3}\textup{-CUT}$. More precisely, the completeness and soundness correspond to $(57/62-\varepsilon,55/62+\varepsilon)$, and the eventual hardness corresponds to $55/57+\varepsilon$.
@@
(a) Gadget reducing $P_{0}^{=}$ to $P^{\neq}$. The vertices $\{1,2,3\}$ are primary variables, and the rest are auxiliary variables. All edges represent single copies of $P^{\neq}$ clauses. The global variables are not pictured as they are unused.
(b) Gadget reducing $P_{1}^{=}$ to $P^{\neq}$. The vertices $\{1,2,3\}$ are primary variables, $\{3,4,5\}$ are global variables, and the rest are auxiliary variables. The edges represent one, two, and three copies of the corresponding $P^{\neq}$ clause, depending on the thickness.
Figure 2: Gadgets found by AlphaEvolve for reducing $3\textup{LIN}({3})$ to $\textup{MAX-}{3}\textup{-CUT}$ (see Appendix C for a more explicit description via edge lists). Note that an edge $(i,j)$ in the graph corresponds to the predicate $P^{\neq}$ applied to variables $i$ and $j$, with thickness proportional to its number of copies in the instance.
Figure 3: Gadget found by AlphaEvolve for reducing $3\textup{LIN}({4})$ to $\textup{MAX-}{4}\textup{-CUT}$. The thickness of an edge $(i,j)$ is proportional to the number of copies of the predicate $P^{\neq}$ applied to variables $i$ and $j$ in the gadget, which is between $1$ and $1429$ (see Appendix C for a more explicit description via edge lists).
@@
To the best of our knowledge, the current best hardness factor for $\textup{MAX-}{4}\textup{-CUT}$ is $85/86+\varepsilon$ [3, Theorem 1.2], which we improve with Theorem 3.2 to $0.987$. The hardness in [3] was obtained by a reduction from $\textup{MAX-}{3}\textup{-CUT}$ to $\textup{MAX-}{k}\textup{-CUT}$, for all $k>3$.

### v6

@@ 1 Introduction
    Hardness of certifying upper bounds on the MAX-CUT and MAX-IND-SET of sparse random graphs. In Section 2, we improve upon the combinatorial structures presented in [42], achieving improved lower bounds in the $\{3,4\}$-regular cases. Complementing this result, we obtain new upper bounds which improve upon Hoffman’s classic spectral bounds [35, 30]. (These upper bounds are derived via analytical approaches [^1] These improvements are derived from a discussion with Sidhanth Mohanty, Northwestern University..) Together, these achieve an almost tight resolution to these problems. [^2] By this we mean: the lower bounds and upper bounds on efficient certification algorithms match up to the second decimal place..
@@
    NP-Hardness of approximation for metric TSP. In Section 4, we give a new gadget based reduction (from standard Håstad PCPs [31]) for the NP-hardness of approximating metric TSP. We improve the current SOTA of $117/116$ [15] to $111/110$. Our improvement advances the long line of work establishing hardness results for this problem [51, $1+\delta$] (without an explicit value for $\delta$), [20, 5381/53805381/5380][10, 3183/31823183/3182][50, 220/219220/219][43, 185/184185/184][39, 123/122123/122][15, $117/116$]. As an aside, our AI-based exploration is ongoing, with the potential for further improvements to this bound.
Notably, our results rely on a single (AI-derived) technique, AlphaEvolve [48, 54], to find and verify finite combinatorial structures that improve on prior constructions. At a high level AlphaEvolve uses an LLM (Large-Language-Model) to iteratively evolve code snippets that generate combinatorial structures (we will sometimes refer to these as constructions) that are of high quality by some criterion. While the structures we generate using AlphaEvolve are finite constructions, our main results (Theorems 2.13.1, and 4.1) embed the universal quantifier $\forall n$, with $n$ being the size of the problem instance, thanks to appropriate “lifting” [42, 59, 15] arguments. At a minimum our work suggests that we would do well to routinely run gadget proofs through an “AI optimization” phase. We present our results in increasing order of human involvement (required to make the problem amenable to AlphaEvolve).
@@
For the hardness of certifying upper bounds for sparse random graphs, we use AlphaEvolve to search for $d=\{3,4\}$-regular Ramanujan graphs, along with a witness for a large cut/independent set. Evaluating whether a graph is Ramanujan, and whether or not the cut/independent set is a valid witness, are efficient. Consequently, for these problems we could find and verify graphs with as many as $163$ vertices.
@@
We emphasize that our theorems do not depend on the correctness of these fast verifiers; even though our AlphaEvolve computations internally use them, the gadgets found by AlphaEvolve were verified by a brute-force algorithm that explicitly checks all $\Omega(k^{m})$ constraints.
2 Hardness of Certification Problems on Random Graphs
A central problem in average case complexity is that of certification, where our goal is to certify a property of typical samples from a distribution. We focus on the setting of sparse random graphs, where the distribution in question is $\mathcal{G}(n,d)$, the uniform distribution over all $n$-vertex $d$-regular graphs, where $d\in\mathbb{N}$ is small. Let $\Omega$ be the set of $n$-vertex graphs, and consider a property $P:\Omega\to\{0,1\}$ such that $\mathop{\mathbf{Pr}}_{G\sim\mathcal{G}(n,d)}[P(G)=1]=1-o_{n}(1)$. We are given as input a graph $G\sim\mathcal{G}(n,d)$, and our algorithmic task is to efficiently certify that $G$ satisfies $P$ with high probability.
@@ 2 Hardness of Certification Problems on Random Graphs
Here, $\mathcal{G}(n,d)$ is the uniform distribution over all $n$-vertex $d$-regular graphs for some fixed constant $d\in\mathbb{N}$.
We focus specifically on the setting of refutation, where $P$ is a co-NP property (recall that a problem is in co-NP if the NO instance of a decision problem has an efficiently verifiable certificate). This setting has seen considerable activity, leading to innovation in various algorithmic techniques such as spectral and sum-of-squares algorithms, along with the development of a variety of methods to establish hardness (for instance, [11, 55, 8, 42, 41]).
In this work, we consider the problem of certifying upper bounds on the MAX-CUT (or MAX-IND-SET) of a graph drawn from $\mathcal{G}(n,d)$. For example, for MAX-CUT, we study the following question: for what values of $\sigma$ can we efficiently certify with high probability that a random graph drawn from $\mathcal{G}(n,d)$ has MAX-CUT at most $\sigma\times\#(\texttt{of edges})$?
@@
There is evidence that this is a nontrivial problem exhibiting a certification gap, in the sense that $\textup{MC}(G)$ concentrates around a constant value strictly smaller than $\sigma^{\textup{MC}}_{d}$ for $G\sim\mathcal{G}(n,d)$ [8, 42]. These represent the structurally most elementary refutation-style properties (involving only binary variables and pairwise constraints) that exhibit interesting complexity-theoretic phenomena. In this work, we focus on accurately characterizing the value of $\sigma^{\textup{MC}}_{d}$ and $\sigma^{\textup{IS}}_{d}$. We remark that the asymptotics of these quantities as $d\to\infty$ are already understood (thanks to [8, 42]). In service of a precise understanding of this problem in all regimes, we focus on the least understood case where $d$ is small; specifically we study $d\in\{3,4\}$.
Towards this goal, [42] introduced a conjecture [^3] We refer the reader to [42] for supporting evidence and intuition. that relates $\sigma^{\textup{MC}}_{d}$ and $\sigma^{\textup{IS}}_{d}$ to a fundamental question in spectral graph theory, specifically about Ramanujan graphs (Definition 2.4 below), which are (in a spectral sense) the most “random” looking regular graphs constructed via a deterministic process.
The intuition behind this framework relies on the strategy of ”quiet planting”. Specifically, if one can exhibit a Ramanujan graph failing to satisfy a certain property $P$ (e.g., a small MAX-CUT value), then under the hardness assumption of [42, Conjecture 1.6], it is impossible for a polynomial-time algorithm to certify $P$ on the random graph distribution $\mathcal{G}(n,d)$. In the context of MAX-CUT, if we construct a Ramanujan graph with a maximum cut of at least $\gamma^{\textup{MC}}_{d}\times\#(\texttt{edges})$, then the certification bound $\sigma^{\textup{MC}}_{d}$ must essentially be at least $\gamma^{\textup{MC}}_{d}$, where $\gamma^{\textup{MC}}_{d}$ is defined in Definition 2.5 below. Therefore, the objective in establishing stronger lower bounds for certification is to maximize this value $\gamma^{\textup{MC}}_{d}$ over the set of Ramanujan graphs. We employ AlphaEvolve to construct such extremal Ramanujan graphs for both MAX-CUT and MAX-IND-SET. The precise conjecture we operate with is stated in 2.1.
@@
[42] exhibited finite $d$-regular Ramanujan graphs for $d\in\{3,4\}$, yielding concrete lower bounds on $\sigma_{d}^{\textup{MC}}$ and $\sigma_{d}^{\textup{IS}}$ that are strong enough to prove a certification gap. On the other hand, there was a large gap between the best known upper bound [35, 30] and their lower bounds (see Table 1 for a comparison). Our results make progress towards a more precise understanding of this question. Firstly, we use AlphaEvolve to construct better $d$-regular Ramanujan graphs, exhibiting stronger lower bounds.
@@ 3 Gadget based NP-Hardness for Approximating MAX-k-CUT
Consider a collection of predicates $\mathcal{P}=\{\Gamma_{1},\ldots,\Gamma_{r}\}$ over some finite alphabet $\Omega$, where the predicates are functions $\Gamma_{i}:\Omega^{l_{i}}\to\{0,1\}$. A CSP is defined by a set of variables and a collection of clauses, where each clause applies a predicate from $\mathcal{P}$ to a subset of the variables, and the goal is to find an assignment of values from $\Omega$ to the variables that maximizes the number of satisfied constraints. The hardness of approximating a CSP is parameterized by two parameters $c$ (completeness) and $s$ (soundness). In this paper we will prove statements of the following form: Assuming $\texttt{P}\neq\texttt{NP}$, it is hard to distinguish between CSP instances where the maximum fraction of satisfied clauses is at least $c$, from those where it is at most $s$. This in turn implies that it is NP-hard to approximate the maximum number of satisfied clauses in a CSP instance within a factor of $s/c$. In this paper we will parameterize the hardness of CSPs with the tuple $(c,s)$.
@@
Let $\Omega$ be a finite alphabet, and let $\Gamma_{1},\ldots,\Gamma_{r}$ be a collection of predicates over $\Omega$, that is, $\Gamma_{i}:\Omega^{\ell_{i}}\to\{0,1\}$ for some $\ell_{i}\in\mathbb{N}$. We define $\textup{CSP}(\Gamma_{1},\ldots,\Gamma_{r})$ to be the collection of instances $\texttt{I}:\Omega^{n}\to\mathbb{R}_{\geq 0}$ of the form $\texttt{I}(\mathbf{x})=\sum_{j\in[m]}\zeta_{j}(\mathbf{x})$ where for each $j$, $\zeta_{j}$ is a clause, i.e., it takes the form $\zeta_{j}(\mathbf{x})=\Gamma_{i}(x_{a_{1}},x_{a_{2}},\ldots,x_{a_{\ell_{i}}})$ for some $i\in[r]$ and distinct $a_{1},\ldots,a_{\ell_{i}}\in[n]$.
In the Max-CSP setting, the input is a description of an instance $I$ of a particular CSP, and the goal is to compute the maximum number of constraints that can be satisfied. In other words, we are trying to approximately compute the function $f(\texttt{I})=\max_{\mathbf{x}\in\Omega^{n}}\texttt{I}(\mathbf{x})$. It will be useful to cast the approximation problem as a decision problem, as we do below by introducing the notion of $(c,s)$-approximation.
@@
Given a constraint satisfaction problem $\textup{CSP}(\Gamma_{1},\ldots,\Gamma_{r})$ and constants $0\leq s\leq c\leq 1$, the $(c,s)$ approximation problem for $\textup{CSP}(\Gamma_{1},\ldots,\Gamma_{r})$ is the problem of, given an $n$-variable instance $\texttt{I}=\sum_{j\in[m]}\zeta_{j}$, deciding whether (1) $\max_{\mathbf{x}}\texttt{I}(\mathbf{x})\geq c\cdot m$ or (2) $\max_{\mathbf{x}}\texttt{I}(\mathbf{x})\leq s\cdot m$. Similarly for $0<\alpha\leq 1$, the $\alpha$-approximation problem is the problem of finding some $\mathbf{x}^{\prime}$ such that $\texttt{I}(\mathbf{x}^{\prime})\geq\alpha\cdot\max_{x}\texttt{I}(\mathbf{x})$.
@@
We have a strong quantitative understanding for a number of CSPs [31, 16] under NP-hardness. However, there are some quite simple CSPs for which very little is known quantitatively. For example, consider $\textup{MAX-}{k}\textup{-CUT}$, the problem of partitioning the vertices of a graph into $k$ sets in a way that maximizes the number of edges crossing the partition. This problem can be written as a CSP over alphabet $\Omega=\mathbb{Z}_{k}$ as $\textup{MAX-}{k}\textup{-CUT}:=\textup{CSP}(\texttt{P}^{\neq}_{2})$ where $\texttt{P}^{\neq}_{2}$ is the “inequality” predicate $\texttt{P}^{\neq}_{2}(x,y)=\mathbf{1}\{x\neq y\}$. For small values of $k$, there is a large gap between the best known algorithmic results (concretely, $0.836$ and $0.857$-approximations for $k=3$ and $k=4$ respectively [28, 18]) and the best known NP-hardness ($16/17+\varepsilon\approx 0.941$ and $85/86+\varepsilon\approx 0.9883$ for $k=3$ and $k=4$ respectively [5]) [^4] Even under the Unique Games Conjecture, the optimal approximability curve is not known for $k=4$, and for $k=3$ is only known in the low completeness regime [32] (which implies a weak inapproximability of $0.989$)..
@@
As an illustrative example, consider the reduction from 3-SAT to MAX-2-SAT appearing in [59]. Here, the source constraint is a 3-CNF clause $C=x_{1}\lor x_{2}\lor x_{3}$. This single constraint is mapped to a gadget consisting of a collection of ten 2-CNF clauses. The local assignment is defined over the original variables $\{x_{1},x_{2},x_{3}\}$ along with an auxiliary variable $z$ introduced specifically for this gadget.
The target collection contains the singleton clauses $x_{1},x_{2},x_{3},z$, the pairwise negated clauses of the variables, and mixed clauses linking the variables to the auxiliary. The full set of target clauses is:
$\begin{gathered}x_{1},x_{2},x_{3},\neg x_{1}\lor\neg x_{2},\neg x_{2}\lor\neg x_{3},\neg x_{3}\lor\neg x_{1},\\ z,\neg x_{1}\lor\neg z,\neg x_{2}\lor\neg z,\neg x_{3}\lor\neg z.\end{gathered}$
In this construction, if the source constraint is satisfied, there exists a local assignment (a value for $z$) such that a fraction $c=0.7$ (7 out of 10) of the target clauses are satisfied. Conversely, if the source constraint is not satisfied, no local assignment can satisfy more than a fraction $s=0.6$ of the target constraints.
We use AlphaEvolve to find such gadget reductions from the well-understood CSP $3\textup{LIN}({k})$ to $\textup{MAX-}{k}\textup{-CUT}$ for $k\in\{3,4\}$. Here $3\textup{LIN}({k})$ is defined as $\textup{CSP}\left(\texttt{P}^{\equiv 0}_{3,k},\texttt{P}^{\equiv 1}_{3,k},\ldots,\texttt{P}^{\equiv(k-1)}_{3,k}\right)$ for the $3$-ary “linear equation” predicates $\texttt{P}^{\equiv i}_{3,k}(x,y,z)=\mathbf{1}\{x+y+z\equiv i\pmod{k}\}$. We adopt the notation $\texttt{P}^{\texttt{type}}_{r,k}$, where the superscript specifies the nature of the constraint, $r$ denotes the arity, and $k$ (if present) indicates the alphabet size.
@@
For any $\varepsilon>0$, there is a gadget reduction from $(1-O(\varepsilon),1/3+O(\varepsilon))$-approximating $3\textup{LIN}({3})$ to $(0.9193-\varepsilon,0.887+\varepsilon)$-approximating $\textup{MAX-}{3}\textup{-CUT}$. As a consequence, it is NP-hard to $0.9649+\varepsilon$-approximate $\textup{MAX-}{3}\textup{-CUT}$. More precisely, the completeness and soundness correspond to $(57/62-\varepsilon,55/62+\varepsilon)$, and the corresponding to a hardness of $55/57+\varepsilon$.
@@
(a) Gadget reducing $\texttt{P}^{\equiv 0}_{3,3}$ to $\texttt{P}^{\neq}_{2}$. The vertices $\{1,2,3\}$ are primary variables, and the rest are auxiliary variables. All edges represent single copies of $\texttt{P}^{\neq}_{2}$ clauses. The global variables are not pictured as they are unused.
(b) Gadget reducing $\texttt{P}^{\equiv 1}_{3,3}$ to $\texttt{P}^{\neq}_{2}$. The vertices $\{1,2,3\}$ are primary variables, $\{3,4,5\}$ are global variables, and the rest are auxiliary variables. The edges represent one, two, and three copies of the corresponding $\texttt{P}^{\neq}_{2}$ clause, depending on the thickness.
Figure 2: Gadgets found by AlphaEvolve for reducing $3\textup{LIN}({3})$ to $\textup{MAX-}{3}\textup{-CUT}$ (see Appendix C for a more explicit description via edge lists). Note that an edge $(i,j)$ in the graph corresponds to the predicate $\texttt{P}^{\neq}_{2}$ applied to variables $i$ and $j$, with thickness proportional to its number of copies in the instance.
Figure 3: Gadget found by AlphaEvolve for reducing $3\textup{LIN}({4})$ to $\textup{MAX-}{4}\textup{-CUT}$. The thickness of an edge $(i,j)$ is proportional to the number of copies of the predicate $\texttt{P}^{\neq}_{2}$ applied to variables $i$ and $j$ in the gadget, which is between $1$ and $1429$ (see Appendix C for a more explicit description via edge lists).
@@
To the best of our knowledge, the current best hardness factor for $\textup{MAX-}{4}\textup{-CUT}$ is $85/86+\varepsilon$ [5, Theorem 1.2], which we improve with Theorem 3.2 to $0.987$. The hardness in [5] was obtained by a reduction from $\textup{MAX-}{3}\textup{-CUT}$ to $\textup{MAX-}{k}\textup{-CUT}$, for all $k>3$.